OpenStudy (anonymous):

Establish the identity: 1/1-(cos^2 theta)/(1+sin theta) = sin theta

zepdrix (zepdrix):

mmmm that's difficult to read. Lemme know if this is formatted correctly,\[\Large\rm \frac{\left(\frac{1}{1-\cos^2\theta}\right)}{1+\sin \theta}=\sin \theta\]Like that..?

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{1}-\cfrac{cos^2(\theta)}{1+sin(\theta)}=sin(\theta)\quad ?\)

OpenStudy (anonymous):

Sorry about that, I'm fairly new to this site. It's \[1- [(\cos \theta ^{2})/(1+\sin \theta)] = \sin \theta\]

OpenStudy (anonymous):

I hope that's somewhat better.

OpenStudy (anonymous):

So you had it comepletely correct.

zepdrix (zepdrix):

Mmm ok here is one approach :) \[\Large\rm 1-\frac{\cos^2\theta}{1+\sin \theta}\color{royalblue}{\left(\frac{1-\sin \theta}{1-\sin \theta}\right)}\] We could multiply this second term, top and bottom, by the conjugate of our denominator.

zepdrix (zepdrix):

The reason we do this is because it will allow us to make use of our Pythagorean Identity. Leave the top alone, multiplying out the bottom gives us:\[\Large\rm 1-\frac{\cos^2\theta(1-\sin \theta)}{1-\sin^2 \theta}\]Does the multiplication make sense?

OpenStudy (anonymous):

Okay, that's a much better approach than mine. The multiplication makes much more sense. Thank you.

zepdrix (zepdrix):

Oh you got it from there? ^^

OpenStudy (anonymous):

I believe so. Thank you so much!!

OpenStudy (anonymous):

\[1-\dfrac{cos^2(x)}{1+sin(x)}=\dfrac{1+sinx}{1+sinx}-\dfrac{cos^2(x)}{1+sin(x)}\] replace \(cos^2 = 1 - sin^2\) \[\dfrac{1+sinx-1+sin^2x}{1+sinx}=\dfrac{sin^2x+sinx}{1+sinx}=\dfrac{sinx(1+sinx)}{1+sinx}=sinx\]

OpenStudy (anonymous):

Thank you, Ooops!

OpenStudy (anonymous):

:)

OpenStudy (kanwal32):

I will tell u a easier way in 2steps